# Primordial black holes from slow phase transitions: a model-building perspective

### v1

@@ Abstract
We investigate the formation of primordial black holes (PBHs) through delayed vacuum decay during slow cosmic first-order phase transitions. Two specific models, the polynomial potential and the real singlet extension of the Standard Model, are used as illustrative examples. Our findings reveal that models with zero-temperature scalar potential barriers are conducive to the realization of this mechanism, as the phase transition duration is extended by the U-shaped Euclidean action. We find that the resulting PBH density is highly sensitive to the barrier height, with abundant PBH formation observed for sufficiently high barriers. Notably, the phase transition needs NOT to be super-strong, and the commonly used exponential nucleation approximation $\Gamma(t)\sim e^{\beta t}$ fails to capture the PBH formation dynamics in such models.
@@ 1 Introduction
Due to their profound cosmological and astrophysical implications, primordial black holes (PBHs) have attracted significant research attention Carr:2020gox ; Green:2020jor ; Escriva:2022duf ; LISACosmologyWorkingGroup:2023njw . Those hypothetical black holes are from the collapse of overdense regions right after the Big Bang, long before the existence of stars and galaxies. Cosmic first-order phase transitions (FOPTs) provide a favorable environment for PBH formation: the Universe transitions from a higher energy false vacuum to a lower energy true vacuum via tunneling, resulting in the true vacuum bubbles nucleate and expand in the false vacuum background. This may lead to the formation of PBHs, if the bubble collisions compress a large amount of energy into a small volume Hawking:1982ga ; Lewicki:2019gmv ; Jung:2021mku ; Balaji:2024rvo , or if some false vacuum remnants form significant overdensities via delayed vacuum decay Kodama:1982sf ; Hall:1989hr ; Kusenko:2020pcg ; Liu:2021svg ; Hashino:2021qoq ; Hashino:2022tcs ; He:2022amv ; Kawana:2022olo ; Lewicki:2023ioy ; Gouttenoire:2023naa ; Gouttenoire:2023pxh ; Salvio:2023blb ; Conaci:2024tlc ; Banerjee:2023brn ; Banerjee:2023vst ; Banerjee:2023qya ; Baldes:2023rqv ; Lewicki:2024ghw ; Jinno:2023vnr ; Flores:2024lng or trapping fermions Lee:1986tr ; Gross:2021qgx ; Baker:2021nyl ; Baker:2021sno ; Kawana:2021tde ; Huang:2022him ; Kawana:2022lba ; Lu:2022jnp ; Marfatia:2021hcp ; Marfatia:2022jiz ; Tseng:2022jta ; Xie:2023cwi ; Acuna:2023bkm ; Lewicki:2023mik ; Chen:2023oew ; Kim:2023ixo ; Gehrman:2023qjn ; DelGrosso:2024wmy ; Borah:2024lml . While FOPTs might be generic during the thermal history of the Universe Mazumdar:2018dfl , the formation of PBHs during such transitions is highly nontrivial. It typically imposes additional constraints on the temperature, strength, duration, and bubble expansion velocity of the FOPT, as well as the particle content of the underlying model. Depending on scenarios, the PBHs from FOPTs have various mass, spin, and abundance distributions.
@@ 2.2 The exponential approximation and beyond
Once the underlying particle model is built, one is able to derive $S(t)$ and also $\Gamma(t)$, and proceed the calculation as described in Section 2.1. To simplify the calculation, in many cases people expand the action at some moment $t_{*}$ as
@@
The other is the percolation time $t_{p}$, the moment that the bubbles in the normal patch forms an infinite connecting cluster, i.e. $F(t_{p})=0.71$. In general $t_{p}>t_{n}$, and when the FOPT is prompt ($\beta/H_{*}\gg 1$) and moderate ($\alpha\lesssim 1$), $t_{p}\approx t_{n}$. However, if the FOPT is very strong or lasts for a long time, $t_{p}$ differs significantly from $t_{n}$. In that case, it is proposed to adopt $t_{*}=t_{p}$, and replace $\beta/H_{*}$ with $(8\pi)^{1/3}v_{w}/(H_{*}\bar{R})$ in the GW calculation, where $\bar{R}$ is the mean separation of bubbles Ellis:2018mja ; Ellis:2019oqb ; Wang:2020jrd , which can be adopted as $[n_{b}(t_{*})]^{-1/3}$, with
@@
the bubble number density in the normal patches. Furthermore, one should also check the FOPT completion condition
@@
such that the physical volume of the false vacuum is decreasing at percolation Turner:1992tz ; Ellis:2018mja .
The exponential approximation is widely used in FOPT research and is generally effective. It is also commonly employed in the study of PBH formation from delayed vacuum decay Liu:2021svg ; Hashino:2021qoq ; Hashino:2022tcs ; He:2022amv ; Kawana:2022olo ; Gouttenoire:2023naa ; Gouttenoire:2023pxh ; Salvio:2023blb ; Conaci:2024tlc ; Banerjee:2023brn ; Banerjee:2023vst ; Banerjee:2023qya ; Baldes:2023rqv ; Lewicki:2024ghw . However, PBH formation predominantly occurs in the small $\beta/H_{*}$ region, which corresponds to slow FOPTs with prolonged durations. For instance, previous studies indicate $\beta/H_{*}\lesssim 10$ for abundant PBH formation Liu:2021svg ; Lewicki:2023ioy ; Gouttenoire:2023naa . When the first-order expansion coefficient $\beta$ of $S(t)$ is small, it becomes necessary to check if truncating the expansion Eq. (14) at linear order provides a satisfactory approximation. This can be determined by examining whether the ratio of the second-order expansion to the first-order one, $\left|[\zeta^{2}(t-t_{*})^{2}/2]/[\beta(t-t_{*})]\right|$, is much less than 1 during the FOPT. This is equivalent to checking
@@ 4 The $\mathbb{Z}_{2}$-symmetric singlet extension of the SM
To provide a comprehensive overview of the parameter space, we conducted a 2-dimensional scan for $m_{s}$ and $\lambda_{hs}$ while keeping $\lambda_{s}$ as different fixed values, as depicted in the left panel of Fig. 8. We specifically plotted the parameter space for $f_{\rm pbh}\in[10^{-100},1]$ and observed that it consists of narrow regions resembling “lines”. These “lines” have endpoints determined by ensuring the triviality bound Lindner:1985uk . It is well known that large scalar coupling constants are in general required to realize the FOPT. However, such large scalar coupling constants are strongly constrained by the triviality bound Kanemura:2012hr . We take into account the constraint from the triviality bound by employing mass dependent beta functions Kanemura:2023wap . We require that the Landau pole of the scalar couplings remains above 10 TeV to satisfy the triviality bound.
The PBH formation parameter space corresponds to strong GW production, and we have evaluated the corresponding GW spectra and verified that they are easily detected by the future LISA LISA:2017pwj , TianQin TianQin:2015yph , Taiji Hu:2017mde , and DECIGO Kawamura:2011zz interferometers. Moreover, we illustrate the deviations in the $hhh$ and $hWW/hZZ$ couplings as purple and green dashed contours, respectively. These deviations can be precisely measured at future collider experiments like HL-LHC Cepeda:2019klc , CEPC An:2018dwb , ILC Bambade:2019fyw , and FCC-ee FCC:2018evy . Those future collider experiments can offer complementary and correlated probes to this PBH formation mechanism.
In the middle and right panels of Fig. 8, we present the variations of $f_{\rm pbh}$ and $m_{\rm pbh}$ with respect to $\lambda_{hs}$, considering different treatments for $S(t)$ while keeping $m_{s}=218$ GeV and $\lambda_{s}=2$ fixed. We observe that the PBH fraction exhibits rapid changes as a function of $\lambda_{hs}$, primarily due to the significant influence of the barrier height on the probability of delayed vacuum decay. Conversely, the PBH mass $\sim 10^{28}\,{\rm g}\,(\sim 10^{-5}M_{\odot})$ changes gradually as it is predominantly determined by the temperature of the FOPT. It can be seen in the figure that the linear expansion fails to accurately describe the dynamics of PBH formation, as expected.
Refer to captionRefer to caption
Figure 8: Left: the parameter space for PBH formation in the $\mathbb{Z}_{2}$-symmetric xSM. Middle (Right): The PBH fraction (mass) as a function of $\lambda_{hs}$ for different treatments of $S(t)$, where $m_{s}=218$ GeV and $\lambda_{s}=2$ are fixed.

### v2

@@ Abstract
We investigate the formation of primordial black holes (PBHs) through delayed vacuum decay during slow cosmic first-order phase transitions. Two specific models, the polynomial potential and the real singlet extension of the Standard Model, are used as illustrative examples. Our findings reveal that models with zero-temperature scalar potential barriers are conducive to the realization of this mechanism, as the phase transition duration is extended by the U-shaped Euclidean action. We find that the resulting PBH density is highly sensitive to the barrier height, with abundant PBH formation observed for sufficiently high barriers. Notably, the phase transition needs not to be ultra-supercooled (i.e. the parameter $\alpha\gg 1$), and the commonly used exponential nucleation approximation $\Gamma(t)\sim e^{\beta t}$ fails to capture the PBH formation dynamics in such models.
@@ 1 Introduction
Due to their profound cosmological and astrophysical implications, primordial black holes (PBHs) have attracted significant research attention Carr:2020gox ; Green:2020jor ; Escriva:2022duf ; LISACosmologyWorkingGroup:2023njw . Those hypothetical black holes originate in the early Universe shortly after the Big Bang, long before the existence of stars and galaxies. Cosmic first-order phase transitions (FOPTs) provide a favorable environment for PBH formation: the Universe transitions from a higher energy false vacuum to a lower energy true vacuum via tunneling, resulting in the true vacuum bubbles nucleate and expand in the false vacuum background. This may lead to the formation of PBHs, if the bubble collisions compress a large amount of energy into a small volume Hawking:1982ga ; Lewicki:2019gmv ; Jung:2021mku , or if some false vacuum remnants form significant overdensities via delayed vacuum decay Kodama:1982sf ; Hall:1989hr ; Kusenko:2020pcg ; Liu:2021svg ; Hashino:2021qoq ; Hashino:2022tcs ; He:2022amv ; Kawana:2022olo ; Lewicki:2023ioy ; Gouttenoire:2023naa ; Gouttenoire:2023pxh ; Salvio:2023blb ; Conaci:2024tlc ; Banerjee:2023brn ; Banerjee:2023vst ; Banerjee:2023qya ; Baldes:2023rqv ; Lewicki:2024ghw ; Balaji:2024rvo ; Jinno:2023vnr ; Flores:2024lng or trapping fermions Lee:1986tr ; Gross:2021qgx ; Baker:2021nyl ; Baker:2021sno ; Kawana:2021tde ; Huang:2022him ; Kawana:2022lba ; Lu:2022jnp ; Marfatia:2021hcp ; Marfatia:2022jiz ; Tseng:2022jta ; Xie:2023cwi ; Acuna:2023bkm ; Lewicki:2023mik ; Chen:2023oew ; Kim:2023ixo ; Gehrman:2023qjn ; DelGrosso:2024wmy ; Borah:2024lml . While FOPTs might be generic during the thermal history of the Universe Mazumdar:2018dfl , the formation of PBHs during such transitions is highly nontrivial. It typically imposes additional constraints on the temperature, strength, duration, and bubble expansion velocity of the FOPT, as well as the particle content of the underlying model. Depending on scenarios, the PBHs from FOPTs have various mass, spin, and abundance distributions.
@@ 2.2 The exponential approximation and beyond
Once the underlying particle model is built, one is able to derive $S(t)$ and also $\Gamma(t)$, and proceed the calculation as described in Section 2.1. To simplify the calculation, most previous studies have chosen to expand the action at a specific moment $t_{*}$,
@@
The other is the percolation time $t_{p}$, which is the moment when the bubbles in the normal patch form an infinite connecting cluster, i.e., $F(t_{p})=0.71$ rintoul1997precise . In general $t_{p}>t_{n}$, and when the FOPT is prompt ($\beta/H_{*}\gg 1$) and moderate ($\alpha\lesssim 1$), $t_{p}\approx t_{n}$. However, if the FOPT is very strong or lasts for a long time, $t_{p}$ differs significantly from $t_{n}$. In that case, it is proposed to adopt $t_{*}=t_{p}$, and replace $\beta/H_{*}$ with $(8\pi)^{1/3}v_{w}/(H_{*}\bar{R})$ in the GW calculation, where $\bar{R}$ is the mean separation of bubbles Ellis:2018mja ; Ellis:2019oqb ; Wang:2020jrd , which can be adopted as $[n_{b}(t_{*})]^{-1/3}$, with
@@
the bubble number density in the normal patches. Furthermore, one should also check the FOPT completion condition Turner:1992tz ; Ellis:2018mja
@@
such that the physical volume of the false vacuum is decreasing at percolation.
The exponential approximation is widely used in FOPT research and is generally effective. It is also commonly employed in the study of PBH formation from delayed vacuum decay Liu:2021svg ; Hashino:2021qoq ; Hashino:2022tcs ; He:2022amv ; Kawana:2022olo ; Gouttenoire:2023naa ; Gouttenoire:2023pxh ; Salvio:2023blb ; Conaci:2024tlc ; Banerjee:2023brn ; Banerjee:2023vst ; Banerjee:2023qya ; Baldes:2023rqv ; Lewicki:2024ghw ; Balaji:2024rvo . However, PBH formation predominantly occurs in the small $\beta/H_{*}$ region, which corresponds to slow FOPTs with prolonged durations. For instance, previous studies indicate $\beta/H_{*}\lesssim 10$ for abundant PBH formation Liu:2021svg ; Lewicki:2023ioy ; Gouttenoire:2023naa . When the first-order expansion coefficient $\beta$ of $S(t)$ is small, it becomes necessary to check if truncating the expansion Eq. (14) at linear order provides a satisfactory approximation. This can be determined by examining whether the ratio of the second-order expansion to the first-order one, $\left|[\zeta^{2}(t-t_{*})^{2}/2]/[\beta(t-t_{*})]\right|$, is much less than 1 during the FOPT. This is equivalent to checking
@@ 4 The $\mathbb{Z}_{2}$-symmetric singlet extension of the SM
To provide a comprehensive overview of the parameter space, we conducted a 2-dimensional scan for $m_{s}$ and $\lambda_{hs}$ while keeping $\lambda_{s}$ as different fixed values, as depicted in the top panel of Fig. 8. In the top-left panel, we specifically plotted the parameter space for $f_{\rm pbh}\in[10^{-100},1]$ and observed that it consists of narrow regions resembling “lines”. These “lines” have endpoints determined by ensuring the triviality bound Lindner:1985uk . It is well known that large scalar coupling constants are in general required to realize the FOPT. However, such large scalar coupling constants are strongly constrained by the triviality bound Kanemura:2012hr . We take into account the constraint from the triviality bound by employing mass dependent beta functions Kanemura:2023wap . We require that the Landau pole of the scalar couplings remains above 10 TeV to satisfy the triviality bound.
The top-right panel of Fig. 8 shows the details of the parameter space for $\lambda_{s}=2$, near the vicinity of $m_{s}=218.4$ GeV and $\lambda_{hs}=1.106$. The PBH formation parameter space corresponds to strong GW production, and we have evaluated the corresponding GW spectra and verified that they are easily detected by the future LISA LISA:2017pwj , TianQin TianQin:2015yph , Taiji Hu:2017mde , and DECIGO Kawamura:2011zz interferometers. Moreover, we illustrate the deviations in the $hhh$ and $hWW/hZZ$ couplings as purple and green dashed contours, respectively. These deviations can be precisely measured at future collider experiments like HL-LHC Cepeda:2019klc , CEPC An:2018dwb , ILC Bambade:2019fyw , and FCC-ee FCC:2018evy . Those future collider experiments can offer complementary and correlated probes to this PBH formation mechanism.
In the bottom-left and -right panels of Fig. 8, we present the variations of $f_{\rm pbh}$ and $m_{\rm pbh}$ with respect to $\lambda_{hs}$, considering different treatments for $S(t)$ while keeping $m_{s}=218$ GeV and $\lambda_{s}=2$ fixed. We observe that the PBH fraction exhibits rapid changes as a function of $\lambda_{hs}$, primarily due to the significant influence of the barrier height on the probability of delayed vacuum decay. Conversely, the PBH mass $\sim 10^{28}\,{\rm g}\,(\sim 10^{-5}M_{\odot})$ changes gradually as it is predominantly determined by the temperature of the FOPT. It can be seen in the figure that the linear expansion fails to accurately describe the dynamics of PBH formation, as expected.
Refer to captionRefer to captionRefer to caption
Figure 8: Top: the parameter space for PBH formation in the $\mathbb{Z}_{2}$-symmetric xSM. Bottom-left (-right): The PBH fraction (mass) as a function of $\lambda_{hs}$ for different treatments of $S(t)$, where $m_{s}=218$ GeV and $\lambda_{s}=2$ are fixed.